Home > Reports > 2005

Report Archive

2017 - 2016
2015 - 2014 - 2013 - 2012
2011 - 2010 - 2009 - 2008
2007 - 2006 - 2005 - 2004

Night Reports 2005

December 2005
Hours spent Observing: 132.0, Weather: 187.6, Technical downtime: 17.7

DateHours Used Comment
ObservingWeatherTechnical
31 December 20059.60.01.3Non-photometric. Leap second requiring system reconfiguration.
30 December 200510.90.00.0Full night.
29 December 20050.010.90.0Ice
28 December 20050.010.90.0High humidity and Ice
27 December 20050.010.90.0High humidity and Ice
26 December 200510.90.00.0 Full night.
25 December 20052.28.70.0High humidity for much of night.
24 December 20055.05.90.0High humidity from midnight.
23 December 20052.48.50.0High humidity for much of night.
22 December 20055.04.91.0High humidity from 1am. Autoguider problems.
21 December 20050.07.03.9High humidity. Mirror cover jammed.
20 December 20050.010.90.0High humidity.
19 December 20050.010.90.0High humidity.
18 December 20050.07.73.2Instrument control computer locked up. High humidity.
17 December 20050.010.90.0High humidity
16 December 20050.010.90.0High humidity
15 December 20050.010.90.0Freezing Fog
14 December 20056.04.40.5High humidity until midnight. Cass node lost, autorecovered.
13 December 200510.30.00.6Mirror cover problems at start of night.
12 December 20050.010.90.0Thick cloud
11 December 20057.43.50.0Thick cloud until 23:00. Non-photometric. Poor seeing.
10 December 20056.94.00.0Thick cloud and high winds from 02:30
09 December 20057.93.00.0Thick cloud until 11:30. Non-Photometric.
08 December 20056.50.04.4 TCS-RCS comms problems
07 December 20058.01.51.4High humidity at start of night. 7x Acc unresponsive autorecovered. Mirror cover problems
06 December 20051.98.80.1Non-photometric. 6x Acc unresponsive autorecovered
05 December 20050.010.80.0Humidity
04 December 200510.60.00.2Non photometric. Acc unresponsive, autorecovered twice.
03 December 20059.80.01.0Non photometric. Acc unresponsive, autorecovered twice. AZM unresponsive, autorecovered.
02 December 200510.70.00.1Non photometric. Acc unresponsive, autorecovered twice.
01 December 20050.010.80.0Thick Cloud

November 2005
Hours spent Observing: 140.0, Weather: 147.3, Technical downtime: 18.4

Reliability and Downtime: Observing and poor weather have been balanced about 50:50 which is disappointing, but it is Winter. Technical downtime has again fallen to 10% after the small increase in the preceding month. We are very pleased with this consistent performance since the mid Summer.

Tropical Storm Delta hit the Canaries at the end of the month causing a lot of damage, very little of which was at the observatory. Our anemometer does not read above 65mph, but the NOT weather station appears to have registered winds exceeding 90mph. Damage to LT included the loss of a sign and a small external out-house, loosening of the weather station rain gauge and damp inside the webcam, so nothing which prevented observations once the weather cleared.

Engineering: There have been three visits to site since the last report. One was predominantly electrical and IT related, working on increasing system robustness and replacing failed hard drives.

The second included TTL engineers and mostly involved software upgrades. The full software infrastructure of high and low-level software (excluding just the TCS) has been updated to the latest TTL release versions. It has been over a year since the last upgrade. Despite a small problem with the focus mechanism which is fully resolved, no observational time was lost due to the software change which was performed in a staged manner over several days. Some minimal work was also possible to investigate the optical astigmatism and the results of that work are still under analysis.

The third visit was for instrument commissioning. Two new instruments, the RINGO polarimeter and the Meaburn Spectrograph have seen first light and are currently undergoing robotic commissioning to allow their routine use. Neither are common user instruments, but further details can be found in a news item on the LT web pages. The bad news on the instrumentation front is the damage to the cryogenic system for SupIRCam. We have already directly contacted all users of this instrument with further details of the problem.

Queen's Anniversary Prize: It has been announced that the work of the ARI, including the Liverpool Telescope has led to Liverpool John Moores University being awarded the biennial Queen's Anniversary Prize for Higher and Further Education, one of the UK's most prestigious educational awards. Further details can be found in a news item on the LT web pages.

DateHours Used Comment
ObservingWeatherTechnical
30 November 20055.54.01.3ACAD encoder errors (1.2). Acc unresponsive, autorecovered (0.1). Clouds drifting through site.
29 November 20050.010.80.0Wind & Rain (Tropical storm Delta) all night. Opportunity taken to vacuum pump RATCam overnight.
28 November 20050.010.80.0Wind & Rain (Tropical storm Delta) all night.
27 November 20050.010.70.0Wind & Rain (Tropical storm Delta) all night. 8xAcc unresponsive, autorecovered.
26 November 20050.010.70.0Thick cirrus. 3xAcc unresponsive, autorecovered.
25 November 20055.25.50.0Thick cirrus. 2xAcc unresponsive, autorecovered. Alt unresponsive, autorecovered
24 November 200510.20.00.4Acc unresponsive. Autorecovered
23 November 20050.710.00.0Non-photometric
22 November 20050.38.22.0Acc disk errors
21 November 2005
20 November 20058.91.00.7High humidity until 20:00, then robotic operations. AZM unreponsive at 00:00, autorecovered ok.
19 November 20054.06.60.0Instrument commisssioning until 10:30pm, then high humidity.
18 November 20056.04.60.0High humidity until 10pm. Instrument commissioning until 2am, then robotic ops (photometric). High humidity from 4am.
17 November 200510.60.00.0Non photometric. Instrument commissioning until 2am, then robotic ops.
16 November 20050.010.60.0Humidity
15 November 20050.010.60.0Humidity
14 November 20050.010.60.0Humidity
13 November 20050.010.60.0Humidity
12 November 20051.98.60.0Humidity & High winds until 4am
11 November 20050.010.50.0Humidity
10 November 200510.50.00.0Photometric.
09 November 20059.00.01.5Non-photometric. Cas unresponsive, autorecovered (0.5). TCS network interface timeouts (1.0)
08 November 20059.50.01.0Non-photometric. Azm unresponsive, autorecovered (1.0).
07 November 20058.10.02.3Non-photometric. Azm unresponsive twice (1.0). TCS network interface time-outs (1.3).
06 November 200510.40.00.0Photometric
05 November 20059.60.00.8ACADS fault.
04 November 200510.40.00.0Full night.
03 November 20055.23.12.1Non-Photometric. High humidity. Focus Problems (1.0), Network Problems (0.6), RCS Problems (0.6).
02 November 20059.30.01.0Photometric. TCS network interface time-outs (1.0).
01 November 20055.00.05.3Photometric. Oil level fault (1.0). AGD fault (1.0). TCS network interface time-outs (1.0). Node disc failure (2.3)

October 2005
Hours spent Observing: 126.4, Weather: 141.3, Technical downtime: 39.4


DateHours Used Comment
ObservingWeatherTechnical
31 October 20055.10.05.1Non-photometric. Focus failed (0.3). Oil level fault (4.8)
30 October 20050.09.21.0Replacing failed science-fold deployment motor. Cloud.
29 October 20051.09.20.0Science-fold deployment failed
28 October 20050.00.010.2Computer memory refurb. (5.0) Telescope in local control. (5.2)
27 October 20050.010.20.0Wind. Humidity.
26 October 20050.010.20.0Wind. Humidity.
25 October 20052.54.33.3Fix and test ACC boot problem (1.5). Focus mechanism init failed (1.7). TCC unresponsive and auto-recovered (0.1). High winds (4.3)
24 October 20054.60.05.5Photometric. ACC crash (0.5). TCC unresponsive and not recovered (5.0))
23 October 200510.10.00.0Photometric. Cass stuck during homing.
22 October 20050.49.70.0High humidity
21 October 200510.10.00.0Non-photometric
20 October 20059.50.50.0Non-photometric
19 October 20050.010.00.0High Humidity
18 October 200510.00.00.0Non-photometric.
17 October 20059.20.00.7Non-photometric. CAS unresponsive. Autorecovered ok.
16 October 20050.09.90.0Rain
15 October 20050.09.90.0Rain
14 October 20050.09.90.0Rain
13 October 20059.10.00.7AMN unresponsive. Autorecovered ok.
12 October 20059.80.00.0Autoguider testing then robotic
11 October 20059.80.00.0Autoguider testing then robotic
10 October 20057.42.30.0Autoguider testing then robotic. High Winds.
9 October 20050.09.70.0Rain and wind.
8 October 20050.09.70.0Heavy rain and wind.
7 October 20050.09.70.0Heavy rain and wind.
6 October 20050.09.70.0Humidity and cloud.
5 October 20058.60.01.1OSS reboot failure
4 October 20054.00.05.6Network loss to azm not autorecovered.
3 October 20055.60.04.0RCS errors in new software version
2 October 20059.60.00.0On-sky system checks (4.0) then robotic (5.6)
1 October 2005---Humidity and cloud.

September 2005
Hours spent Observing: 63.8, Weather: 66.9, Technical downtime: 1.3

Reliabilty and Downtime: A significant amount of time was lost this month as a consequence of the brush fires on La Palma (this is noted as "Weather" in the statistics). We would like to thank all the employees of the ORM and staff from several other telescopes whose efforts saved the observatory site facilities from serious damage. From the numbers above, you will see that excluding the scheduled maintenance, the telescope downtime resulting from technical faults has again been very low this month.

Engineering: The September engineering trip went very well with almost everything we had hoped for being completed. The mirror was recoated and new zero points will be published soon on the RATCam information web page. A first look reduction shows a 0.3 mag increase in SDSS r' throughput. Work on the cassegrain rotator appears to have significantly improved its performance, which should in turn bring pointing and tracking improvements. The optics have been recollimated better than they were before the engineering work and investigations are ongoing to see if work on the secondary mechanism has had any effect, beneficial or detrimental on the focussing. A new pointing model has been installed with an RMS pointing error of about 8 arcsec. For the first two weeks of October, a small fraction of each night will be spent on further engineering and performance tests, but the telescope is effectively back in full science operations.

Imager Field Rotation: The most pleasing outcome of the downtime is that the cassegrain de-rotator cable wrap has been reorganised in such a way as to allow a full 360 degree travel. Users can now specify the field rotation they would like for each observation. For all targets currently loaded in the phase II database, a default orientation of 'sky position angle = 0' has been assumed. If you require a field orientation other than this, please contact the support astronomer, Chris Moss, just as you would to make any modification to your observing strategy.

Support: Until now, users have been only been able to access their data through the "Recent Data" pages. You can now download your data from the archive search results screen. You will still need to input your proposal ID and password.

Time Applications: Applications for PATT semester 2006A have closed. For semester 2006A, the PATT oversubscription rate for available time was 3.8. Following a release of engineering time to supplement the available PATT time, the oversubscription rate was readjusted to 1.8 (324.5 hours requested for a total of 180.8 hours available for the semester; in addition, 235.2 hours have been pre-allocated to the semester from PATT proposals in previous semesters.) Calls are expected to be made soon for the JMU TAG and ITP international time. Announcements will be made on the News web page.
 

DateHours Used Comment
ObservingWeatherTechnical
30 September 2005---Scheduled maintenance & Commissioning. Two hours on-sky testing and then rain.
29 September 2005---Scheduled maintenance & Commissioning.
28 September 2005---Scheduled maintenance & Commissioning.
27 September 2005---Scheduled maintenance & Commissioning.
26 September 2005---Scheduled maintenance & Commissioning.
25 September 2005---Scheduled maintenance & Commissioning.
24 September 2005---Scheduled maintenance & Commissioning.
23 September 2005---Scheduled maintenance & Commissioning.
22 September 2005---Scheduled maintenance & Commissioning.
21 September 2005---Scheduled maintenance & Commissioning.
20 September 2005---Scheduled maintenance & Commissioning.
19 September 2005---Scheduled maintenance & Commissioning.
18 September 2005---Scheduled maintenance & Commissioning.
17 September 2005---Scheduled maintenance & Commissioning.
16 September 2005---Scheduled maintenance & Commissioning.
15 September 20059.00.00.0Full robotic night. Photometric.
14 September 20059.00.00.0Full robotic night. Photometric.
13 September 20050.09.00.0High humidity and rain all night.
12 September 20053.65.00.3AMN lost and autorecovred ok. High humidity for 2nd half of night
11 September 20050.08.90.0Fire pre-alert continutes.
10 September 20050.08.90.0Fire pre-alert continutes. As does high humidity.
9 September 20050.08.80.0Fire pre-alert continutes. Also high humidity!
8 September 20050.08.80.0One fire still threatens ORM, no site inspection possible until ORM alert status lifted.
7 September 20050.08.80.0Fires are out, but still too much smoke in the air to observe.
6 September 20050.08.70.0Forest fires. No known damage, but operations suspended until a site inspection is possible.
5 September 20058.70.00.0Non-photometric.
4 September 20058.70.00.0Non-photometric.
3 September 20057.60.01.0Cass stuck during homing, auto recovered (0.0). Network loss to cas auto-recovered (1.0). Non-photometric.
2 September 20058.60.00.0Cass stuck during homing. Auto recovered. Photometric.
1 September 20058.60.00.0Photometric. RATCam only.

August 2005
Hours spent Observing: 173.3, Weather: 63.4, Technical downtime: 13.1

First I would like to thank all those who contributed science results to the recent telescope operations report for PPARC. It was good to see so many interesting results starting to come from the telescope. Equally we thank those who gave feedback about the telescope performance limitations which had prevented you yet producing any results from the LT. All these comments and suggestions have been discussed and we have been able to respond positively to many of you.

Reliability: The most obvious headline this month has to be the extremely encouraging increase in telescope reliability. This has certainly been our best month yet with only 5% technical downtime and only one full night being lost. The weather has been a little disappointing for August.

If you read the nightly reports, you will see a variety of small one-off problems which have been fixed as the come up. The only two re-occurring problems to have lost us time this month have been the mysterious network losses to the various nodes and initialisation errors of the focus mechanism. The latter has been fully resolved in software. The former is still under investigation.

Performance: Those receiving data will know that there continues to be work to be done on the tracking performance of the telescope, however significant improvements have been made since the last report. Following the re-alignment of the Azimuth encoder system at the end of June, the fraction of severely trailed images has dropped from 10% to 2%. There will be an extended period of engineering down-time in September which is to include installation of the new auto-guider to help address the tracking limitations.

We are now also using a new pointing model which was derived following the encoder realignment. Pointing repeatability within a night is better than 10 arcsec RMS and absolute accuracy from night to night adds a further 10 arcsec RMS error.

Support: There have been further minor changes in the data archive and MyLT web pages. Both have been made more robust and the "Recent Data" pages at Recent Data have been subdivided by proposal ID. We hope this makes it easier for you to quickly locate your data.

The application guidelines have been fully revised for PATT semester 06A. Please take a look before submitting your application.

There are still a few programmes which have been allocated time in semester 2005B for which we have not received phase 2 observation specification forms. As soon as we get these forms, the data can be loaded to the database and be ready for observation. Of course, if you have target-of-opportunity time, you obviously cannot submit forms early.


DateHours Used Comment
ObservingWeatherTechnical
31 August 20058.50.00.0Photometric. RATCam only.
30 August 20058.50.00.0Photometric. RATCam only.
29 August 20050.00.08.5Cass stuck during homing. Auto recovered. Unrelated simultaneous failure of both instruments.
28 August 20058.40.00.0Cass stuck during homing. Auto recovered. Photometric
27 August 20058.30.00.1Network loss to tcc auto-recovered. Non-photometric
26 August 20050.08.40.0Cloud.
25 August 20050.08.30.0Thick Cloud.
24 August 20058.30.00.0Non-photometric.
23 August 20058.30.00.0Non-photometric.
22 August 20058.20.00.0Non-photometric.
21 August 20050.08.20.0Thick cloud. High humidity.
20 August 20050.08.20.0Thick cloud.
19 August 20050.08.20.0Humidity. Thick cloud.
18 August 20050.08.10.0Humidity
17 August 20052.16.00.0Pointing test at start of night, then cloud and humidity
16 August 20058.10.00.0Photometric
15 August 20058.10.00.0Pointing test first half of night.
14 August 20057.90.00.1PLC node 36 error on power-up. Site visit to fix.
13 August 20050.08.00.0Cloud
12 August 20057.90.00.0Non-photometric.
11 August 20055.60.02.3Network loss to csc auto-recovered (0.5). Computer disk error (1.8). Photometric.
10 August 20057.90.00.0Photometric.
9 August 20057.80.00.0Photometric.
8 August 20057.00.00.8Focus mechanism failed to initialise correctly. Photometric.
7 August 20057.00.00.8Cass stuck during homing. Auto recovered. Focus mechanism failed to initialise correctly (0.8). Photometric.
6 August 20057.70.00.0Cass stuck during homing. Auto recovered.
5 August 20057.70.00.0Cass stuck after ROTEX. Photometric
4 August 20057.20.00.5Focus mechanism failed to initialise correctly. Maybe photometric
3 August 20057.60.00.0Photometric
2 August 20057.60.00.0Photometric
1 August 20057.60.00.0Photometric

July 2005
Hours spent Observing: 131.1, Weather: 51.9, Technical downtime: 40.4

DateHours Used Comment
ObservingWeatherTechnical
31 July 20057.10.30.2Altitude instability (0.2). High humidity at end of night (0.3).
30 July 20056.80.00.7Azimuth became unresponsive. Auto recovered OK.
29 July 20050.07.50.0High humidity all night.
28 July 20050.07.40.0High humidity all night.
27 July 20050.07.40.0High humidity all night. Opportunity taken to pump SupIRCam.
26 July 20050.07.40.0High humidity all night
25 July 20056.40.01.0SupIRCam temperature problems
24 July 20056.30.01.1TCC unresponsive, auto reboot (0.1). Failure of TCS network interface commands (1.0)
23 July 20054.00.03.3Large (several arcmin) pointing offsets (1.5). Failure of TCS network interface commands (1.5)
22 July 20050.07.30.0Calima much reduced on previous nights
21 July 20050.07.30.0Calima
20 July 20050.07.30.0Calima (i.e. dust)
19 July 20050.00.07.2Scheduled engineering downtime. CCD vacuum pumping
18 July 20050.00.07.2Focus mechanism failed to initialise correctly.
17 July 20057.20.00.0Full Robotic night. Photometric.
16 July 20056.20.01.0Photometric. SupIRCam computer manually rebooted
15 July 20057.10.00.0Full robotic night. Photometric.
14 July 20056.50.00.6Robotic night. Photometric. Altitude became unresponsive. Autorecovered OK
13 July 20057.10.00.0Full robotic night. Photometric
12 July 20056.10.01.0NAS faults. Photometric
11 July 20056.50.00.6AZM lost homing. Autorecovered OK. Photometric
10 July 20057.10.00.0Full robotic night. Photometric
9 July 20054.00.03.0Photometric
8 July 20056.00.01.0Mix of manual & robotic. Photometric. Auto reboot of tcc time lock failure.
7 July 20057.00.00.0Full night of observing. Mix of manual & robotic. Photometric.
6 July 20057.00.00.0Full night of observing. Mix of manual & robotic. Photometric.
5 July 20057.00.00.0Full night of observing. Mix of manual & robotic. Photometric.
4 July 20050.00.07.0Oil Level Fault.
3 July 20051.50.05.5Secondary mirror problems
2 July 20057.00.00.0Full robotic night
1 July 20056.90.00.0Full robotic night

June 2005
Hours spent Observing: 108.3, Weather: 8.8, Technical downtime: 97.3

Another mixed month in terms of hours observed. The first half of the month was very good, but a lot of time has been lost in the second half of the month to the failure of the mirror cover to open. This has now been fixed. Whilst on the subject of telescope problems, users should note that several times this month the telescope was going in and out of focus during the night. The badly defocussed data has not been distributed to users.

Engineering: The promised site visit by a joint TTL / ARI team has taken place. A full review of the performance of the hydraulic system was performed, but it is too early to say if there will be any major changes in telescope performance. Some work was also done on the Azimuth encoders which has made a very significant improvement the telescope tracking.

Support: Dr. Chris Moss has been making contact with many observers and discussion with them has started to increase observing efficiency for several programmes through increasing our understanding of user requirements. In more tangible changes, there have been some modifications to the data archive. Most of these will not be obvious to the user, but there have been a few formatting changes to the archive search tools (Data Archive). Most significantly, the new data subdirectories available at Recent Data are now arranged to put all the data from a given night together rather than the previous practice of filing them according to UT date which led to each night's data getting split across two subdirectories.

Email Contact: Finally, we wanted to make sure everyone was aware that emails get sent to project PIs to notify them as soon as any data has been obtained and are available in the archive. Several PIs have not been receiving these emails due to the configuration of firewalls and email SPAM filters. If you know that data have been obtained for your programme, but you have not had an email about them, please let us know so that we can monitor the extent of this problem. You can always check the new data page (Recent Data) or MyLiverpoolTelescope to see if you have any new data available.

DateHours Used Comment
ObservingWeatherTechnical
30 June 20053.90.03.0Pointing tests following hydraulic system adjustments
29 June 20055.90.01.0After completion of mirror cover work (1.0), no faults all night.
28 June 20050.00.06.9Operations suspended awaiting mirror cover work. Work commenced, but incomplete.
27 June 20050.00.06.9Operations suspended awaiting mirror cover work
26 June 20050.00.06.9Operations suspended awaiting mirror cover work
25 June 20050.00.06.9Operations suspended awaiting mirror cover work
24 June 20051.40.05.5Jammed mirror cover makes robotic operation impossible. Maintenance staff travelling out to site on Monday. Some manual pointing and image quality data collected.
23 June 20050.00.06.9Jammed mirror cover.
22 June 20050.00.06.9Jammed mirror cover.
21 June 20056.90.00.0Full robotic night (photometric)

20 June 2005

4.0

0.0

2.8

Network loss to Az not auto-recovered.
19 June 20050.00.06.8Jammed mirror cover.
18 June 20050.00.06.8Jammed mirror cover.
17 June 20050.05.41.5Heavy cloud. Jammed mirror cover.
16 June 20050.03.43.5Heavy cloud. Jammed mirror cover.
15 June 20056.80.00.1Network loss to tcc auto-recovered. Non-photometric.
14 June 20056.90.00.0Full robotic night. Photometric.
13 June 20054.00.02.9Network loss to tcc auto-recovered (0.3). Focus mechanism instability (2.6). Photometric.
12 June 20050.00.06.9A&G box circuit breaker tripped on start-up.
11 June 20054.60.02.3Network loss to cas auto-recovered (0.5). Instrument control computer hang (1.8). Photometric.
10 June 20056.60.00.3Network loss to tcc auto-recovered. Photometric.
09 June 20050.00.06.9Defocussed by 0.4mm all night. No data. Under investigation.
08 June 20057.00.00.0Full robotic night. Not Photometric.
07 June 20057.00.00.0Full robotic night. Photometric.
06 June 20057.00.00.0Full robotic night. Photometric.
05 June 20057.00.00.0Full robotic night. Photometric.
04 June 20052.70.04.3Network loss to Az not auto-recovered. Photometric.
03 June 20057.00.00.0Full robotic night. Photometric.
02 June 20055.70.01.3Series of node losses/reboots autorecovered. Photometric.
01 June 20057.00.00.0Full robotic night. Photometric.

May 2005
Hours spent Observing: 124.4, Weather: 27.2, Technical downtime: 79.2

Science: We have had a real science highlight this month, with our best yet fully automated GRB response. The LT obtained multi-colour photometry of GRB050502 starting 3 minutes after the Integral gamma-ray detection. See the news story on the LT News page or in the news archive.

Engineering: Telescope reliabilty is up again this month after a disappointing April. Following my upbeat comments in the April report, the first week of May proved very poor. The two main problems were a sticky Cass bearing and a hydraulic pump motor failure. The latter was replaced and has not given us any further problems. Since this is the second pump replaced this year (see Jan 2005), a full recheck of the hydraulic scavenge system has been scheduled for June. The cass bearing problems are currently more mysterious and TTL will be helping us investigate this also in June. In the meantime, we have a work-around which has kept us observing most of the month.
We spent a couple of nights on a more thorough analysis of the pointing performance. The great majority of pointings are currently within 30arcsec of the field centre, though there still occaissional unexplained errors up to 2 arcmin.

Support: It has become clear that the differences between the way the LT and traditional observatories operate has caused a significant stumbling block for many observers preventing optimum exploitation of the telescope. This month, Dr Chris Moss has joined the LT operations group as our new support astronomer. He will not be involved in any routine maintenance and operations, allowing him to concentrate on helping all users obtain the maximum scientific returns from their observing time. He will be overseeing and providing expert advice throughout the application, scheduling and observing process. Revisions to the submission procedures will be announced soon, but in the mean time he can be contacted via the LT Helpdesk at this address (updated 5th Nov 2007).

Other: The Live Data page has been made public online. You can check on the LT weather station and the state of various telescope mechanisms at any time. On moonlit nights you can even see the telescope moving on the webcam.

DateHours Used Comment
ObservingWeatherTechnical
31 May 20057.10.00.0Full robotic night. Very little cirrus.
30 May 20057.10.00.0Full robotic night. Photometric
29 May 20050.00.07.1Enclosure PLC PSU failure
28 May 20057.10.00.0Full robotic night. Little cloud
27 May 20057.10.00.0Full robotic night. Little cloud
26 May 20051.55.70.0High humidity
25 May 20057.20.00.0Pointing model work first half of night
24 May 20056.30.01.0Pointing model work first half of night. 1 hour network loss to AZM auto-recovered.
23 May 20057.30.00.0Full night of robotic operation
22 May 20050.00.07.3PLC fault preventing AMN boot
21 May 20050.00.07.3PLC fault preventing AMN boot
20 May 20050.00.07.3PLC fault preventing AMN boot
19 May 20057.30.00.0Full night of robotic operation
18 May 20056.90.00.5Focus tests. 0.5hrs lost to rotator instability.
17 May 20056.00.01.4cloud (non-photometric). Network loss auto-recovered
16 May 20052.44.01.0Lots of cloud. Network loss to AMN auto-recovered
15 May 20050.07.40.0Thick Cloud
14 May 20050.07.50.0Thick Cloud
13 May 20053.50.04.0Oil Level Fault
12 May 20057.50.00.0Fully robotic night. (not photometric)
11 May 20057.50.00.0Fully robotic night. (potentially photometric)
10 May 20057.60.00.0Fully robotic night. (potentially photometric)
09 May 20053.52.61.5High winds. Network loss to Az node auto-recovered.
08 May 20050.00.07.6Cassegrain problems. Experimental workaround installed following day.
07 May 20057.70.00.0Spare pump fitted. Full night of operations
06 May 20050.00.07.7Hydraulic scavenge pump offline.
05 May 20050.00.07.8Hydraulic pump still offline. Parts obtained.
04 May 20050.00.07.8Cass problem cleared. Hydraulic pump still offline.
03 May 20050.00.07.9Cass bearing errors. Sticking? Hydraulic pump failure.
02 May 20057.90.00.0Full night.
01 May 20055.90.02.0

April 2005
Hours spent Observing: 78.7, Weather: 73.2, Technical downtime: 93.4

Engineering: The first half of the night was blighted by degrading Azimuth bearing oil flow performance. A site visit was therefore scheduled for mid-month. Following the work on 20/05/05, no further time has been lost and the telescope has performed reliably for the last week of the month.

DateHours Used Comment
ObservingWeatherTechnical
30 Apr 20050.07.90.0High humidity
29 Apr 20058.00.00.0Full night of robotic data

27 Apr 2005

8.0

0.0

0.0

Full night of robotic data (photometric)

26 Apr 2005

6.0

0.0

2.0

Network problem

25 Apr 2005

0.0

8.1

0.0

Thick clouds

24 Apr 2005

8.1

0.0

0.0

Cirrus (non-photometric). Observed all night without problems

23 Apr 2005

0.0

8.1

0.0

Thick Cloud

22 Apr 2005

8.1

0.0

0.0

Thin cirrus. New pointing model derived and applied. Science Operations for rest of night. No failures.

21 Apr 2005

0.0

8.2

0.0

Thick clouds

20 Apr 2005

0.0

8.2

0.0

Azimuth guide pad adjustment carried out. Telescope available again for operations. Thick clouds on site however.

19 Apr 2005

0.0

0.0

8.2

Operations suspended awaiting site inspection of azimuth

18 Apr 2005

0.0

0.0

8.2

Azimuth problems

17 Apr 2005

6.0

0.0

2.3

Azimuth problems at end of night (photometric)

16 Apr 2005

0.0

8.4

0.0

Freezing Fog

15 Apr 2005

0.0

8.4

0.0

Freezing Fog

14 Apr 2005

5.6

2.1

0.8

Azimuth axis interlock + Auxiliary Mechanism Node issue, both automatically recovered. High humidity at end of night

13 Apr 2005

0.0

0.0

8.5

Azimuth problems

12 Apr 2005

0.0

8.6

0.0

Thick Cloud

11 Apr 2005

0.0

0.0

8.6

Azimuth problems

10 Apr 2005

0.0

0.0

8.6

Azimuth problems

9 Apr 2005

2.0

0.0

6.6

Azimuth problems

8 Apr 2005

5.6

3.0

0.0

High humidity after 03:30

7 Apr 2005

5.6

1.0

2.1

High humidity, air compressor stopped before dawn

6 Apr 2005

6.0

1.2

1.5

Azimuth problem

5 Apr 2005

1.3

0.0

7.5

Azimuth problems

4 Apr 2005

0.5

0.0

8.3

Azimuth guide bearing low flow warning, azimuth homing problems

3 Apr 2005

0.0

0.0

8.8

Cassegrain rotator stuck; axis torque error during homing

2 Apr 2005

0.0

0.0

8.8

Cassegrain rotator stuck; axis torque error during homing

1 Apr 20057.90.00.9Azimuth homing problems

March 2005
Hours spent Observing: 121, Weather: 132, Technical downtime: 38.
First half of month spent under the tail end of the blizzards that prevented observing during February. Observations were back underway around mid-March.

DateHours Used Comment
ObservingWeatherTechnical

31 Mar 2005

4.9

4.0

0

High humidity

30 Mar 2005

0.3

0.0

8.6

Azimuth homing

29 Mar 2005

8.0

0.0

1.0

Azimuth lost homing three times during the night

28 Mar 20058.60.00.4Azimuth lost homing, autorecovered ok, (non-photometric)
27 Mar 20059.10.00.0Full night of robotic data (photometric)
26 Mar 20059.10.00.0Full night of robotic data (photometric)
25 Mar 20059.10.00.0Full night of robotic data (photometric)
24 Mar 20056.23.00.0High humidity until 23:30, then non-photometric
23 Mar 20059.20.00.0Full night of robotc data (photometric)
22 Mar 20058.50.00.7Altitude Node Failed, Autorecovery worked
21 Mar 20050.09.20.0High humidity
20 Mar 20050.70.08.6Problems with wetness sensor
19 Mar 20056.33.00.0Thick cloud at start of night, non-photometric
18 Mar 20057.70.01.6Some initial focus problems, a lot of cloud, non-photometric
17 Mar 20058.00.01.4Azimuth lost homing at start of night. Recovered ok. (non-photometric)
16 Mar 20059.40.00.0Full Night of robotic data (non-photometric)
15 Mar 20059.40.00.0Full Night of robotic data (non-photometric)
14 Mar 20052.00.07.4Aux Mech Watchdog / CANBus failure
13 Mar 20050.09.40High humidity
12 Mar 20051.08.50High humidity
11 Mar 20051.08.50High humidity
10 Mar 20051.008.5Air Compressor Fault
09 Mar 20051.08.60
08 Mar 200509.60Weather clear. Enclosure hydraulic mechanism still clogged with ice.
07 Mar 200509.60Weather clear. Enclosure hydraulic mechanism still clogged with ice.
06 Mar 200509.70Weather clear. Enclosure hydraulic mechanism still clogged with ice.
05 Mar 200509.70
04 Mar 200509.70
03 Mar 200509.80SPLC code reloaded.
02 Mar 200509.80SPLC failure. Time would have been lost to this if it were not for the weather which is still foul.
01 Mar 200509.80

February 2005
Time used (Observing:Weather:Fault) 0 : 283.9 : 0

Weather: Incredible as it may seem, we have not opened the enclosure once in February, entirely due to weather. We do not yet have archival LT weather data on-line, but you can see how bad it has been from the ING weather station. For two weeks of the month, the telescope was not contactable due to a circuit breaker supplying the IT equipment tripping and it being impossible to gain access to site to reset it. Two LT staff made an emergency visit to help Dirk with the anticipated clean-up. In fact things were not as bad as feared though there is still significant water ingress to the enclosure during high winds. The new top-seal guttering has made a great difference. The telescope has been started up and run under local supervision. We expect it to be available as soon as the weather clears and ice melts. The autoguider has also been returned to Birkenhead and TTL are working on fixing the problems we have had. Of course none of the work done during the January maintenance visit has yet been properly tested on-sky.

Automation: On the automation front, general system robustness testing has continued despite being unable to observe and we have removed the restriction which stopped the telescope re-opening during the night after any technical fault (see Dec. 2004 report). This means that any minor problems which can be fixed by the automated fault recovery routines will no longer lead to the loss of the rest of the night.

Science: In science/data matters, the downtime has not been wasted. We have taken the opportunity to finally put on-line the new web pages (http://telescope.livjm.ac.uk/), the searchable data archive and phase-II database access under the catchy title "My Liverpool Telescope".

DateHours Used Comment
ObservingWeatherTechnical
28 Feb 200509.80Freezing fog. Everything outdoors cased in 2-3" of ice.
27 Feb 200509.80Fog and gales. Heavy rain
26 Feb 200509.90(non-freezing) Fog and gales.
25 Feb 200509.90Cleared for a few hours hours. Then freezing fog. WHT did manage to open. Seems we may be slightly more prone to ice in the enclosure mechanism than them. Experience long term will tell.
24 Feb 200509.90Guess what. It's snowing.
23 Feb 2005010.00Still ice and 100% humidity.
22 Feb 2005010.00Still ice and 100% humidity.
21 Feb 2005010.00Gained access to LT and powered up. Looks good so far.
20 Feb 2005010.00LT site closed.
19 Feb 2005010.00LT site closed.
18 Feb 2005010.10Road cleared but it started snowing again. Due to insufficient infrastructure (e.g., medical backup), no attempt will be made to access site over the coming weekend even in the unlikely event of the road clearing.
17 Feb 2005010.10Road clearance proceeding, but reached a difficult section.
16 Feb 2005010.10Road clearance proceeding. LT will certainly still be offline 16/02/2005
15 Feb 2005010.10Road clearance is underway. Prediction is to get to the telescopes on Thursday
14 Feb 2005

0

10.1

0

Snow.
13 Feb 2005

0

10.2

0

Snow. Site evacuated.
12 Feb 2005

0

10.2

0

Snow. It's white.
11 Feb 2005

0

10.2

0

More Snow. Mountain top cut off.
10 Feb 2005

0

10.2

0

More Snow. Mountain top cut off.
09 Feb 2005

0

10.3

0

Snow
08 Feb 2005

0

10.3

0

Ice, fog
07 Feb 2005

0

10.3

0

Ice, fog
06 Feb 2005

0

10.3

0

Ice remains, and high winds pre-alert
05 Feb 2005

0

10.4

0

Ice returned and prevented opening enclosure
04 Feb 2005

0

10.4

0

Ice melted during day, freezing fog at night
03 Feb 2005

0

10.4

0

Ice, fog
02 Feb 2005

0

10.4

0

May be clearing. Still too much ice on enclosure to open.
01 Feb 2005

0

10.4

0

Ice.

January 2005
Time used (Observing:Weather:Fault) 93.7 : 133.4 : 102.5
This month we had a "IT & mechanical servicing mission" with two staff from ARI joining Dirk on-site for a week. During their visit they installed a variety of new IT equipment, installed new webcams (accessable from this web site), made a long-term fix to the enclosure locks (see Dec. 2004 report) and portal guide rollers, investigated oil flow limit problems (see 11-13/01/05) and ran further tests on ongoing autoguider problems (we now think the chip is not salvageable). Most of the above were resolved successfully and we are anticipating an increased system robustness over the next couple of months. Weather has been terrible as you can see.

On the automation side, we obtained twilight flat-fields using the fully automated routines for the first time. This will hopefully be integrated into the daily autonomous operations in the coming month.

DateHours Used Comment
ObservingWeatherTechnical
31 Jan 2005

0

10.5

0

Ice
30 Jan 2005

0

10.5

0

Ice
29 Jan 2005

0

10.5

0

Ice
28 Jan 2005

0

10.6

0

Lots of ice
27 Jan 2005

0

10.6

0

Lots of ice
26 Jan 2005

0

10.6

0

Freezing fog
25 Jan 2005

4.0

2.0

4.6

01:19. Oil level fault.
24 Jan 2005

10.6

0

0

23 Jan 2005

3.0

0

7.6

22:39:37. mcc lost contact with alt node. No further diagnostics available in messages log.
22 Jan 2005

7.6

0

3.0

21 Jan 2005

10.6

0

0

20 Jan 2005

0

0

10.7

Continuing work on hydraulic scavenge pump.
19 Jan 2005

0

10.7

0

Pre-alerta for storms overnight. Seems weather was not bad in the event.
18 Jan 2005

0

10.7

0

Snow. New scavenge pump motor installed.
17 Jan 2005

0

10.7

0

Snow
16 Jan 2005

0

10.7

0

Snow
15 Jan 2005

0.3

7.0

3.4

14 Jan 2005

10.7

0

0

13 Jan 2005

2.0

0

8.7

Scavenge pump circuit breaker tripped. Majority of night lost.
12 Jan 2005

1.5

0

9.2

Azimuth oil flow tripped again.
11 Jan 2005

1.5

0

9.2

Azimuth oil flow tripped. Robotic software cannot reset hydraulic faults without supervision.
10 Jan 2005

10.8

0

0

Full robotic night of data.
9 Jan 2005

0

0

10.8

8 Jan 2005

4.0

0

6.8

7 Jan 2005

5.5

0

5.3

6 Jan 2005

0

0

10.8

5 Jan 2005

0

2.0

8.8

4 Jan 2005

0

10.8

0

3 Jan 2005

10.8

0

0

Full robotic night of data.
2 Jan 2005

10.8

0

0

Full robotic night of data.
1 Jan 2005

0

10.9

0